OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given points m=2(8,3)

hero (hero):

\[m(x - x_1) = (y - y_1)\]

hero (hero):

\[2(x - 8) = ( y - 3)\]

hero (hero):

\[2x - 16 = y - 3\]

hero (hero):

\[y = 2x -13\]

OpenStudy (anonymous):

thank u